\begin{thebibliography}{10}

\bibitem{key-1}
A.~Bayad and T.~Kim, \emph{Higher recurrences for Apostol-Bernoulli-Euler numbers}, Russ. J. Math. Phys.
\textbf{19} (2012), no.~1, 1--10.

 \bibitem{key-19}
D.~Ding and J.~Yang, \emph{Some identities related to the Apostol-Euler and Apostol-Bernoulli polynomials},
  Adv. Stud. Contemp. Math. (Kyungshang) \textbf{20}
  (2010), no.~1, 7--21.

\bibitem{key-2}
L.-C. Jang, C.~S. Ryoo, J.~J. Seo, and H.~I. Kwon, \emph{Some properties of the
  twisted {C}hanghee polynomials and their zeros}, Appl. Math. Comput.
  \textbf{274} (2016), 169--177. \MR{3433125}

\bibitem{key-3}
D.~S. Kim and T.~Kim, \emph{A note on {B}oole polynomials}, Integral Transforms
  Spec. Funct. \textbf{25} (2014), no.~8, 627--633. \MR{3195946}

\bibitem{key-5}
\bysame, \emph{Some identities for {B}ernoulli numbers of the second kind
  arising from a nonlinear differential equation}, Bull. Korean. Math. Soc.
  \textbf{52} (2015), no.~6, 2001--2010.

\bibitem{key-4}
\bysame, \emph{Some identities of {K}orobov-type polynomials associated with
  {$p$}-adic integrals on {$\Bbb Z_p$}}, Adv. Difference Equ. (2015), 2015:282,
  13. \MR{3395618}

\bibitem{key-7}
D.~S. Kim and T.~Kim, T.~Komatsu, and S.-H. Lee, \emph{Barnes-type {D}aehee of the
  first kind and poly-{C}auchy of the first kind mixed-type polynomials}, Adv.
  Difference Equ. (2014), 2014:140, 22. \MR{3259855}

\bibitem{key-8}
D.~S. Kim, T.~Kim, and J.~J. Seo, \emph{A note on {C}hanghee polynomials and
  numbers}, Adv. Studies Theor. Phys. \textbf{7} (2013), no.~20, 993--1003.

\bibitem{key-10}
\bysame, \emph{A note on {$q$}-analogue of {B}oole polynomials}, Appl. Math.
  Inf. Sci. \textbf{9} (2015), no.~6, 3135--3158. \MR{3386346}

\bibitem{key-6}
T.~Kim, D. S.~Kim, \emph{A note on non-linear {C}hanghee differential equations}, Russ.
  J. Math. Phys. \textbf{23} (2016), no.~1, 1--6.

\bibitem{key-9}
T.~Kim, \emph{Non-{A}rchimedean {$q$}-integrals associated with multiple
  {C}hanghee {$q$}-{B}ernoulli polynomials}, Russ. J. Math. Phys. \textbf{10}
  (2003), no.~1, 91--98. \MR{2013106 (2004h:33034)}

\bibitem{key-11}
T.~Kim, T.~Mansour, S.~H. Rim, and J.~J. Seo, \emph{A note on $q$-{C}hanghee
  polynomials and numbers}, Adv. Studies Theor. Phys. \textbf{8} (2014), no.~1,
  35--41.
\bibitem{key-12}
H. I.~Kwon, T.~Kim and J. J.~Seo, \emph{ A note on degenerate Changhee numbers and polynomials},
Proc. Jangjeon Math. Soc.18 (2015), no. 3, 295--305.

\bibitem{key-13}
D.~Lim and F.~Qi, \emph{On the appell type--changhee polynomials}, J.
  Nonlinear Sci. \textbf{9} (2016), 1872--1876.

\bibitem{key-14}
J.-W. Park, \emph{On the twisted $q$-{C}hanghee polynomials of higher order},
  J. Comput. Anal. Appl. \textbf{20} (2016), no.~1, 424--431.

\bibitem{key-15}
S.-H. Rim, J.-W. Park, S.-S. Pyo, and J.~Kwon, \emph{The $n$-th twisted
  {C}hanghee polynomials and numbers}, Bull. Korean. Math. Soc. \textbf{52}
  (2015), no.~3, 741--749.

\bibitem{key-16}
J.~J. Seo and T.~Kim, \emph{$p$-adic invariant integral on $\mathbb{Z}_p$
  associated with the {C}hanghee's $q$-bernoulli polynomials}, Int. J. Math.
  Anal. (Ruse) \textbf{7} (2013), no.~41-44, 2117--2128.

\bibitem{key-17}
G.~Y. Sohn and J.~K. Kwon, \emph{A note on twisted {C}hanghee polynomials and
  numbers with weight}, Appl. Math. Sci. \textbf{9} (2015), no.~31, 1517--1525.

\bibitem{key-18}
N.~L. Wang and L.~Hailong, \emph{Some identities on the higher-order {D}aehee
  and {C}hanghee numbers}, Pure and Applied Mathematics Journal \textbf{4}
  (2015), no.~5-1, 33--37.

\end{thebibliography}

\documentclass{amsart}
\usepackage{amstext}
\usepackage{amsthm}
\usepackage{amssymb}

\makeatletter
\numberwithin{equation}{section}
\numberwithin{figure}{section}
\theoremstyle{plain}
\newtheorem{thm}{Theorem}
  \theoremstyle{plain}
  \newtheorem{lem}[thm]{Lemma}

\allowdisplaybreaks
\usepackage[sort&compress,square,numbers]{natbib}

\makeatother

\begin{document}

\title{Differential equations associated with $\lambda$-Changhee polynomials }

\author{Taekyun Kim}
\address{Department of Mathematics, Kwangwoon University, Seoul 139-701, Republic
of Korea}
\email{tkkim@kw.ac.kr}

\author{Dae San Kim}
\address{Department of Mathematics, Sogang University, Seoul 121-742, Republic
of Korea}
\email{dskim@sogang.ac.kr}

\begin{abstract}
In this paper, we study linear differential equations arising from
$\lambda$-Changhee polynomials (or called degenerate Changhee polynomials)
and give some explicit and new identities for the $\lambda$-Changhee
polynomials associated with linear differential equations.
\end{abstract}

\thanks{\noindent {\footnotesize{ \it 2010 Mathematics Subject
Classification } : 05A19, 11B37, 34A30}} \medskip
\thanks{\footnotesize{ \bf Key words and phrases}:
$\lambda$-Changhee polynomials, differential equations}

\maketitle
\global\long\def\relphantom#1{\mathrel{\phantom{{#1}}}}
\global\long\def\Ch{\mathrm{Ch}}

\section{Introduction}

For $N\in\mathbb{N}$, we define the generalized harmonic numbers as
follows:

\begin{equation}
H_{N,0}=1,\quad\text{ for all }N,\label{eq:1}
\end{equation}

\begin{equation}
H_{N,1}=H_{N}=1+\frac{1}{2}+\cdots+\frac{1}{N},\quad\left(\text{see \cite{key-5}}\right),\label{eq:2}
\end{equation}

\begin{equation}
H_{N,j}=\frac{H_{N-1,j-1}}{N}+\frac{H_{N-2,j-1}}{N-1}+\cdots+\frac{H_{j-1,j-1}}{j},\quad\left(2\le j\le N\right).\label{eq:3}
\end{equation}

For $k\in\mathbb{N}$ and $N,j\in\mathbb{N}\cup\left\{ 0\right\} $,
we define the generalized Changhee power sums $S_{k,j}\left(N\right)$
as follows:
\begin{align}
S_{k,0}\left(N\right) & =\left(N+1\right)^{k},\label{eq:4}\\
S_{k,j}\left(N\right) & =\sum_{l=0}^{N}S_{k,j-1}\left(l\right),\quad\left(j\ge1\right),\quad\left(\text{see \cite{key-5,key-6}}\right).\label{eq:5}
\end{align}

In particular, for $k=1$, we also define $S_{1,-1}\left(N\right)=1$.

As is well known, the Euler polynomials are defined by the generating
function
\begin{equation}
\frac{2}{e^{t}+1}e^{xt}=\sum_{n=0}^{\infty}E_{n}\left(x\right)\frac{t^{n}}{n!},\quad\left(\text{see \cite{key-1,key-9,key-10,key-11}}\right).\label{eq:6}
\end{equation}

With the viewpoint of deformed Euler polynomials, the Changhee polynomials
are defined by the generating function
\begin{equation}
\frac{2}{t+2}\left(t+1\right)^{x}=\sum_{n=0}^{\infty}\Ch_{n}\left(x\right)\frac{t^{n}}{n!},\quad\left(\text{see \cite{key-2,key-3,key-8}}\right).\label{eq:7}
\end{equation}

From (\ref{eq:6}) and (\ref{eq:7}), we note that
\begin{align}
 & \sum_{n=0}^{\infty}E_{n}\left(x\right)\frac{t^{n}}{n!}\label{eq:8}\\
 & =\sum_{m=0}^{\infty}\Ch_{m}\left(x\right)\frac{1}{m!}\left(e^{t}-1\right)^{m}\nonumber \\
 & =\sum_{n=0}^{\infty}\left(\sum_{m=0}^{n}\Ch_{m}\left(x\right)S_{2}\left(n,m\right)\right)\frac{t^{n}}{n!},\nonumber
\end{align}
where $S_{2}\left(n,m\right)$ are the Stirling numbers of the second
kind. Thus, by (\ref{eq:8}), we get
\begin{equation}
E_{n}\left(x\right)=\sum_{m=0}^{n}\Ch_{m}\left(x\right)S_{2}\left(n,m\right)\quad\left(n\ge0\right).\label{eq:9}
\end{equation}

The Stirling numbers of the first kind $S_{1}\left(n,l\right)$ appear in the expansion of the falling factorial
\begin{align}
\left(x\right)_{0} & =1,\quad\left(x\right)_{n}=x\left(x-1\right)\cdots\left(x-n+1\right)\label{eq:10}\\
 & =\sum_{l=0}^{n}S_{1}\left(n,l\right)x^{l},\quad\left(n\ge1\right).\nonumber
\end{align}

From (\ref{eq:10}), we note that the generating function of the Stirling
numbers of the first kind is given by
\begin{equation}
\left(\log\left(1+t\right)\right)^{n}=n!\sum_{m=n}^{\infty}S_{1}\left(m,n\right)\frac{t^{m}}{m!},\quad\left(\text{see \cite{key-4,key-7,key-12}}\right).\label{eq:11}
\end{equation}

By (\ref{eq:2}), we easily get
\begin{align}
 & \sum_{n=0}^{\infty}\Ch_{n}\left(x\right)\frac{t^{n}}{n!}\label{eq:12}\\
 & =\sum_{m=0}^{\infty}E_{m}\left(x\right)\frac{1}{m!}\left(\log\left(1+t\right)\right)^{m}\nonumber \\
 & =\sum_{n=0}^{\infty}\left(\sum_{m=0}^{n}S_{1}\left(n,m\right)E_{m}\left(x\right)\right)\frac{t^{m}}{m!}.\nonumber
\end{align}

Thus, by (\ref{eq:12}), we have
\[
\Ch_{n}\left(x\right)=\sum_{m=0}^{n}S_{1}\left(n,m\right)E_{m}\left(x\right),\quad\left(n\ge0\right),\quad\left(\text{see \cite{key-16,key-17}}\right).
\]

Recently, $\lambda$-Changhee polynomials (or called degenerate Changhee
polynomials) are defined by the generating function
\begin{align}
 & \frac{2\lambda}{2\lambda+\log\left(1+\lambda t\right)}\left(1+\frac{\log\left(1+\lambda t\right)}{\lambda}\right)^{x}\label{eq:14}\\
 & =\sum_{n=0}^{\infty}\Ch_{n,\lambda}\left(x\right)\frac{t^{n}}{n!},\quad\left(\text{see \cite{key-13}}\right).\nonumber
\end{align}

When $x=0$, $\Ch_{n,\lambda}=\Ch_{n,\lambda}\left(0\right)$ are
called $\lambda$-Changhee numbers (or called degenerate Changhee
numbers).

In \cite{key-5}, Kim-Kim gave some explicit and new identities for
the Bernoulli numbers of the second kind arising from nonlinear differential
equations. It is known that some interesting identities and properties
of the Frobenius-Euler polynomials are also derived from the non-linear
differential equations (see \cite{key-6,key-11}).

Recently, several authors have studied some interesting properties
for the Changhee numbers and polynomials (see \cite{key-1,key-2,key-3,key-4,key-5,key-6,key-7,key-8,key-9,key-10,key-11,key-12,key-13,key-14,key-15,key-16,key-17,key-18, key-19}).

In this paper, we develop some new method for obtaining identities
related to $\lambda$-Changhee polynomials arising from linear differential
equations. From our study, we derive some explicit and new identities
for the $\lambda$-Changhee polynomials.

\section{Some identities for the $\lambda$-Changhee polynomials arising from
linear differential equations}

First, we introduce lemma for the generalized Changhee power sum $S_{k,j}\left(N\right)$.
\begin{lem}
\label{lem:1} For $2\le r\le N$ and $1\le i\le r-1$, we have
\begin{equation}
S_{1,i-1}\left(r-1-i\right)+S_{1,i-2}\left(r-i\right)=S_{1,i-1}\left(r-i\right).\tag{{A}}\label{eq:lem-1-a}
\end{equation}
\end{lem}
\begin{proof}
From (\ref{eq:4}) and (\ref{eq:5}), we have
\begin{align}
S_{1,i-1}\left(r-1-i\right)+S_{1,i-2}\left(r-i\right) & =\sum_{l=0}^{r-1-i}S_{1,i-2}\left(l\right)+S_{1,i-2}\left(r-i\right)\label{eq:16}\\
 & =\sum_{l=0}^{r-i}S_{1,i-2}\left(l\right)\nonumber \\
 & =S_{1,i-1}\left(r-i\right).\nonumber \qedhere
\end{align}
\end{proof}
Let
\begin{align}
F & =F\left(t;x,\lambda\right)\label{eq:17}\\
 & =\frac{2\lambda}{2\lambda+\log\left(1+\lambda t\right)}\left(1+\lambda^{-1}\log\left(1+\lambda t\right)\right)^{x}.\nonumber
\end{align}

Then, by (\ref{eq:17}), we get
\begin{align}
 & F^{\left(1\right)}\label{eq:18}\\
 & =\frac{d}{dt}F\left(t;x,\lambda\right)\nonumber \\
 & =\lambda\left(1+\lambda t\right)^{-1}\left(-\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-1}+x\left(\lambda+\log\left(1+\lambda t\right)\right)^{-1}\right)F.\nonumber
\end{align}
\begin{align}
 & F^{\left(2\right)}\label{eq:19}\\
 & =\frac{dF^{\left(1\right)}}{dt}\nonumber \\
 & =\lambda^{2}\left(1+\lambda t\right)^{-2}\left\{ \left(2\lambda+\log\left(1+\lambda t\right)\right)^{-1}-x\left(\lambda+\log\left(1+\lambda t\right)\right)^{-1}\right.\nonumber \\
 & \relphantom =+2\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-2}-2x\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-1}\left(\lambda+\log\left(1+\lambda t\right)\right)^{-1}\nonumber \\
 & \relphantom =\left.+\left(x\right)_{2}\left(\lambda+\log\left(1+\lambda t\right)\right)^{-2}\right\} F.\nonumber
\end{align}

So, we are led to put
\begin{align}
 & F^{\left(N\right)}\label{eq:20}\\
 & =\left(\frac{d}{dt}\right)^{N}F\left(t;x,\lambda\right)\nonumber \\
 & =\lambda^{N}\left(1+\lambda t\right)^{-N}\left(\sum_{1\le i+j\le N}a_{i,j}^{\left(\lambda\right)}\left(N,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{-j}\right)F,\nonumber
\end{align}
where $N=1,2,\dots$, and the sum is over all nonnegative integers
$i,j$ with $1\le i+j\le N$.

On the one hand, by (\ref{eq:20}), we get
\begin{align}
 &\relphantom{=} F^{\left(N+1\right)}\label{eq:21}\\
 & =\frac{dF^{\left(N\right)}}{dt}\nonumber \\
 & =\lambda^{N+1}\left(1+\lambda t\right)^{-\left(N+1\right)}\nonumber \\
 & \relphantom =\times\left\{ \left(-N\right)\sum_{\substack{1\le i+j\le N\\i\ge0,j\ge0}}a_{i,j}^{\left(\lambda\right)}\left(N,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{-j}\right.\nonumber \\
 & \relphantom =-\sum_{\substack{2\le i+j\le N+1\\
i\ge1,\,j\ge0
}
}ia_{i-1,j}^{\left(\lambda\right)}\left(N,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{-j}\nonumber \\
 & \relphantom =\left.+\sum_{\substack{2\le i+j\le N+1\\
i\ge0,j\ge1
}
}\left(x-j+1\right)a_{i,j-1}^{\left(\lambda\right)}\left(N,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\nonumber\right.\\
&\relphantom{=}\times\left.\left(\lambda+\log\left(1+\lambda t\right)\right)^{-j}\right\} F.\nonumber
\end{align}

On the other hand, by replacing $N$ by $N+1$ in (\ref{eq:20}),
we have
\begin{align}\label{eq:22}
F^{\left(N+1\right)} & =\lambda^{N+1}\left(1+\lambda t\right)^{-\left(N+1\right)} \\
 & \relphantom =\times\left(\sum_{\substack{1\le i+j\le N+1\\
i,j\ge0
}
}a_{i,j}^{\left(\lambda\right)}\left(N+1,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\right.\nonumber\\
&\relphantom{===}\times\left.\left(\lambda+\log\left(1+\lambda t\right)\right)^{-j}\right)F.\nonumber\\
\end{align}

Let $i+j=r$. Then $1\le r\le N+1$. Comparing the terms with $r=1$,
we get
\begin{align}
a_{1,0}^{\left(\lambda\right)}\left(N+1,x\right) & =-Na_{1,0}^{\left(\lambda\right)}\left(N,x\right)\label{eq:23}\\
a_{0,1}^{\left(\lambda\right)}\left(N+1,x\right) & =-Na_{0,1}^{\left(\lambda\right)}\left(N,x\right).\nonumber
\end{align}

Comparing the terms with $i+j=r$ $\left(2\le r\le N\right)$,
\begin{align}
 & \sum_{i=0}^{r}a_{i,r-i}^{\left(\lambda\right)}\left(N+1,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{i-r}\label{eq:24}\\
 & =-N\sum_{i=0}^{r}a_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{i-r}\nonumber \\
 & \relphantom =-\sum_{i=1}^{r}ia_{i-1,r-i}^{\left(\lambda\right)}\left(N,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{i-r}\nonumber \\
 & \relphantom =+\sum_{i=0}^{r-1}\left(x+i-r+1\right)a_{i,r-i-1}^{\left(\lambda\right)}\left(N,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{i-r}.\nonumber
\end{align}

Thus, by (\ref{eq:24}), we get
\begin{align}
a_{0,r}^{\left(\lambda\right)}\left(N+1,x\right) & =-Na_{0,r}^{\left(\lambda\right)}\left(N,x\right)+\left(x-r+1\right)a_{0,r-1}^{\left(\lambda\right)}\left(N,x\right),\label{eq:25}\\
a_{r,0}^{\left(\lambda\right)}\left(N+1,x\right) & =-Na_{r,0}^{\left(\lambda\right)}\left(N,x\right)-ra_{r-1,0}^{\left(\lambda\right)}\left(N,x\right),\label{eq:26}
\end{align}
and
\begin{align}\label{eq:27}
&\relphantom{=}a_{i,r-i}^{\left(\lambda\right)}\left(N+1,x\right)\\
&=-Na_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)-ia_{i-1,r-i}^{\left(\lambda\right)}\left(N,x\right)\nonumber\\
&\relphantom{=}+\left(x+i-r+1\right)a_{i,r-i-1}^{\left(\lambda\right)}\left(N,x\right),\nonumber
\end{align}
where $1\le i\le r-1$.

Comparing the terms with $i+j=N+1$, we get
\begin{align}
 & \sum_{i=0}^{N+1}a_{i,N+1-i}^{\left(\lambda\right)}\left(N+1,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{i-\left(N+1\right)}\label{eq:28}\\
 & =-\sum_{i=1}^{N+1}ia_{i-1,N+1-i}^{\left(\lambda\right)}\left(N,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{i-\left(N+1\right)}\nonumber \\
 & \relphantom =+\sum_{i=0}^{N}\left(x+i-N\right)a_{i,N-i}^{\left(\lambda\right)}\left(N,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{i-\left(N+1\right)}.\nonumber
\end{align}

From (\ref{eq:28}), we note that
\begin{align}\label{eq:29}
&a_{0,N+1}^{\left(\lambda\right)}\left(N+1,x\right)=\left(x-N\right)a_{0,N}^{\left(\lambda\right)}\left(N,x\right),\\
&a_{N+1,0}^{\left(\lambda\right)}\left(N+1,x\right)=-\left(N+1\right)a_{N,0}^{\left(\lambda\right)}\left(N,x\right),\nonumber
\end{align}
and
\begin{equation}
a_{i,N+1-i}^{\left(\lambda\right)}\left(N+1,x\right)=-ia_{i-1,N+1-i}^{\left(\lambda\right)}\left(N,x\right)+\left(x+i-N\right)a_{i,N-i}^{\left(\lambda\right)}\left(N,x\right),\label{eq:30}
\end{equation}
where $1\le i\le N$.

From (\ref{eq:18}) and (\ref{eq:20}), we have
\begin{align}
 & \lambda\left(1+\lambda t\right)^{-1}\left(-\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-1}+x\left(\lambda+\log\left(1+\lambda t\right)\right)^{-1}\right)F\label{eq:31}\\
 & =F^{\left(1\right)}\nonumber \\
 & =\lambda\left(1+\lambda t\right)^{-1}\left(a_{1,0}^{\left(\lambda\right)}\left(1,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-1}\right.\nonumber\\
 &\relphantom{=}\left.+a_{0,1}^{\left(\lambda\right)}\left(1,x\right)\left(\lambda+\log\left(1+\lambda t\right)\right)^{-1}\right)F.\nonumber
\end{align}

By comparing the coefficients on both sides of (\ref{eq:31}),
we have
\begin{equation}
a_{1,0}^{\left(\lambda\right)}\left(1,x\right)=-1,\quad a_{0,1}^{\left(\lambda\right)}\left(1,x\right)=x.\label{eq:32}
\end{equation}

From (\ref{eq:23}), we note that
\begin{equation}
a_{1,0}^{\left(\lambda\right)}\left(N+1,x\right)=\left(-1\right)^{N+1}N!,\quad a_{0,1}^{\left(\lambda\right)}\left(N+1,x\right)=\left(-1\right)^{N}N!x.\label{eq:33}
\end{equation}

By (\ref{eq:29}), we easily get
\begin{equation}
a_{N+1,0}^{\left(\lambda\right)}\left(N+1,x\right)=\left(-1\right)^{N+1}\left(N+1\right)!,\quad a_{0,N+1}^{\left(\lambda\right)}\left(N+1,x\right)=\left(x\right)_{N+1}.\label{eq:34}
\end{equation}

For $i=1$ in (\ref{eq:30}), we have
\begin{align}
 & a_{1,N}^{\left(\lambda\right)}\left(N+1,x\right)\label{eq:35}\\
 & =-a_{0,N}^{\left(\lambda\right)}\left(N,x\right)+\left(x+1-N\right)a_{1,N-1}^{\left(\lambda\right)}\left(N,x\right)\nonumber \\
 & =-\left(x\right)_{N}+\left(x+1-N\right)a_{1,N-1}^{\left(\lambda\right)}\left(N,x\right)\nonumber \\
 & =-\left(x\right)_{N}+\left(x+1-N\right)\left(-\left(x\right)_{N-1}+\left(x+2-N\right)a_{1,N-2}^{\left(\lambda\right)}\left(N-1,x\right)\right)\nonumber \\
 & =-2\left(x\right)_{N}+\left(x+1-N\right)\left(x+2-N\right)a_{1,N-2}^{\left(\lambda\right)}\left(N-1,x\right)\nonumber \\
 & \vdots\nonumber \\
 & =-N\left(x\right)_{N}+\left(x+1-N\right)\left(x+2-N\right)\cdots\left(x+N-N\right)a_{1,0}^{\left(\lambda\right)}\left(1,x\right)\nonumber \\
 & =-\left(N+1\right)\left(x\right)_{N}\nonumber \\
 & =-S_{1,0}\left(N\right)\left(x\right)_{N}.\nonumber
\end{align}

For $i=2$ in (\ref{eq:30}), we note that
\begin{align}
&\relphantom{=}a_{2,N-1}^{\left(\lambda\right)}\left(N+1,x\right) \label{eq:36}\\
& =-2a_{1,N-1}^{\left(\lambda\right)}\left(N,x\right)\nonumber\\
 & \relphantom{=} +\left(x+2-N\right)a_{2,N-2}^{\left(\lambda\right)}\left(N,x\right)\nonumber \\
 & =\left(-1\right)^{2}2N\left(x\right)_{N-1}\nonumber\\
 &\relphantom{=}+\left(x+2-N\right)\left\{ \left(-1\right)^{2}2\left(N-1\right)\left(x\right)_{N-2}+\left(x+3-N\right)a_{2,N-3}^{\left(\lambda\right)}\left(N-1,x\right)\right\} \nonumber \\
 & =\left(-1\right)^{2}2\left\{ N+\left(N-1\right)\right\} \left(x\right)_{N-1}+\left(x+2-N\right)\left(x+3-N\right)a_{2,N-3}^{\left(\lambda\right)}\left(N-1,x\right)\nonumber \\
 & \vdots\nonumber \\
 & =\left(-1\right)^{2}2\left\{ N+\left(N-1\right)+\left(N-2\right)+\cdots+2\right\} \left(x\right)_{N-1}\nonumber\\
 &\relphantom{=}+\left(x+2-N\right)\left(x+3-N\right)\cdots xa_{2,0}^{\left(\lambda\right)}\left(2,x\right)\nonumber \\
 & =\left(-1\right)^{2}2\left\{ N+\left(N-1\right)+\cdots+2+1\right\} \left(x\right)_{N-1}\nonumber \\
 & =\left(-1\right)^{2}2!S_{1,1}\left(N-1\right)\left(x\right)_{N-1}.\nonumber
\end{align}

Let $i=3$ in (\ref{eq:30}). Then we have
\begin{align}
 & a_{3,N-2}^{\left(\lambda\right)}\left(N+1,x\right)\label{eq:37}\\
 & =-3a_{2,N-2}^{\left(\lambda\right)}\left(N,x\right)+\left(x+3-N\right)a_{3,N-3}^{\left(\lambda\right)}\left(N,x\right)\nonumber \\
 & =\left(-1\right)^{3}3!S_{1,1}\left(N-2\right)\left(x\right)_{N-2}+\left(x+3-N\right)a_{3,N-3}^{\left(\lambda\right)}\left(N,x\right)\nonumber \\
 & \vdots\nonumber \\
 & =\left(-1\right)^{3}3!\left\{ S_{1,1}\left(N-2\right)+\cdots+S_{1,1}\left(1\right)\right\} \left(x\right)_{N-2}\nonumber\\
 &\relphantom{=}+\left(x+3-N\right)\left(x+4-N\right)\cdots xa_{3,0}^{\left(\lambda\right)}\left(3,x\right)\nonumber \\
 & =\left(-1\right)^{3}3!\left\{ S_{1,1}\left(N-2\right)+\cdots+S_{1,1}\left(1\right)+S_{1,1}\left(0\right)\right\} \left(x\right)_{N-2}\nonumber \\
 & =\left(-1\right)^{3}3!S_{1,2}\left(N-2\right)\left(x\right)_{N-2}.\nonumber
\end{align}

Continuing this process, we get
\begin{equation}
a_{i,N+1-i}^{\left(\lambda\right)}\left(N+1,x\right)=\left(-1\right)^{i}i!S_{1,i-1}\left(N-i+1\right)\left(x\right)_{N-i+1},\quad\left(1\le i\le N\right).\label{eq:38}
\end{equation}

Let $2\le r\le N$. Then, by (\ref{eq:25}), we get
\begin{align}
a_{0,r}^{\left(\lambda\right)}\left(N+1,x\right) & =\left(x-r+1\right)a_{0,r-1}^{\left(\lambda\right)}\left(N,x\right)-Na_{0,r}^{\left(\lambda\right)}\left(N,x\right)\label{eq:39}\\
 & =\left(x-r+1\right)a_{0,r-1}^{\left(\lambda\right)}\left(N,x\right)\nonumber\\
 &\relphantom{=}-N\left\{ \left(x-r+1\right)a_{0,r-1}^{\left(\lambda\right)}\left(N-1,x\right)-\left(N-1\right)a_{0,r}^{\left(\lambda\right)}\left(N-1,x\right)\right\} \nonumber \\
 & =\left(x-r+1\right)\left\{ a_{0,r-1}^{\left(\lambda\right)}\left(N,x\right)-Na_{0,r-1}^{\left(\lambda\right)}\left(N-1,x\right)\right\}\nonumber\\
 &\relphantom{=} +\left(-1\right)^{2}N\left(N-1\right)a_{0,r}^{\left(\lambda\right)}\left(N-1,x\right)\nonumber \\
 & \vdots\nonumber \\
 & =\left(x-r+1\right)\sum_{i=0}^{N-r}\left(-1\right)^{i}\left(N\right)_{i}a_{0,r-1}^{\left(\lambda\right)}\left(N-i,x\right)\nonumber\\
 &\relphantom{=}+\left(-1\right)^{N-r+1}N\left(N-1\right)\cdots ra_{0,r}^{\left(\lambda\right)}\left(r,x\right)\nonumber \\
 & =\left(x-r+1\right)\sum_{i=0}^{N-r+1}\left(-1\right)^{i}\left(N\right)_{i}a_{0,r-1}^{\left(\lambda\right)}\left(N-i,x\right).\nonumber
\end{align}

Now, we give an explicit expression for $a_{0,r}^{\left(\lambda\right)}\left(N+1,x\right)$
$\left(2\le r\le N\right)$.

For $r=2$ in (\ref{eq:39}), we have
\begin{align}
a_{0,2}^{\left(\lambda\right)}\left(N+1,x\right) & =\left(x-1\right)\sum_{i=0}^{N-1}\left(-1\right)^{i}\left(N\right)_{i}a_{0,1}^{\left(\lambda\right)}\left(N-i,x\right)\label{eq:40}\\
 & =\left(x\right)_{2}\left(-1\right)^{N-1}\sum_{i=0}^{N-1}\left(N\right)_{i}\left(N-i-1\right)!\nonumber \\
 & =\left(x\right)_{2}\left(-1\right)^{N-1}N!\sum_{i=0}^{N-1}\frac{1}{N-i}\nonumber \\
 & =\left(x\right)_{2}\left(-1\right)^{N-1}N!\sum_{i=1}^{N}\frac{1}{i}\nonumber \\
 & =\left(x\right)_{2}\left(-1\right)^{N-1}N!H_{N,1}.\nonumber
\end{align}

Let us consider $r=3$ in (\ref{eq:39}). From (\ref{eq:40}), we
note that
\begin{align}
a_{0,3}^{\left(\lambda\right)}\left(N+1,x\right) & =\left(x-2\right)\sum_{i=0}^{N-2}\left(-1\right)^{i}\left(N\right)_{i}a_{0,2}^{\left(\lambda\right)}\left(N-i,x\right)\label{eq:41}\\
 & =\left(x-2\right)\sum_{i=0}^{N-2}\left(-1\right)^{i}\left(N\right)_{i}\left(x\right)_{2}\left(-1\right)^{N-i-2}\left(N-i-1\right)!H_{N-i-1}\nonumber \\
 & =\left(x\right)_{3}\left(-1\right)^{N-2}N!\sum_{i=0}^{N-2}\frac{H_{N-i-1}}{N-i}\nonumber \\
 & =\left(x\right)_{3}\left(-1\right)^{N-2}N!H_{N,2}.\nonumber
\end{align}

For $r=4$ in (\ref{eq:39}), we have
\begin{align}
 & a_{0,4}^{\left(\lambda\right)}\left(N+1,x\right)\label{eq:42}\\
 & =\left(x-3\right)\sum_{i=0}^{N-3}\left(-1\right)^{i}\left(N\right)_{i}a_{0,3}^{\left(\lambda\right)}\left(N-i,x\right)\nonumber \\
 & =\left(x-3\right)\sum_{i=0}^{N-3}\left(-1\right)^{i}\left(N\right)_{i}\left(x\right)_{3}\left(-1\right)^{N-i-3}\left(N-i-1\right)!H_{N-i-1,2}\nonumber \\
 & =\left(x\right)_{4}\left(-1\right)^{N-3}N!\sum_{i=0}^{N-3}\frac{H_{N-i-1,2}}{N-i}\nonumber \\
 & =\left(x\right)_{4}\left(-1\right)^{N-3}N!\left\{ \frac{H_{N-1,2}}{N}+\frac{H_{N-2,2}}{N-1}+\cdots+\frac{H_{2,2}}{3}\right\} \nonumber \\
 & =\left(x\right)_{4}\left(-1\right)^{N-3}N!H_{N,3}.\nonumber
\end{align}

Continuing this process, we get
\begin{equation}
a_{0,r}^{\left(\lambda\right)}\left(N+1,x\right)=\left(x\right)_{r}\left(-1\right)^{N-r+1}N!H_{N,r-1}.\label{eq:43}
\end{equation}

For $2\le r\le N$, by (\ref{eq:26}), we get
\begin{align}
 & a_{r,0}^{\left(\lambda\right)}\left(N+1,x\right)\label{eq:44}\\
 & =-ra_{r-1,0}^{\left(\lambda\right)}\left(N,x\right)-Na_{r,0}^{\left(\lambda\right)}\left(N,x\right)\nonumber \\
 & \vdots\nonumber \\
 & =-r\sum_{i=0}^{N-r}\left(-1\right)^{i}\left(N\right)_{i}a_{r-1,0}^{\left(\lambda\right)}\left(N-i,x\right)+\left(-1\right)^{N-r+1}N\left(N-1\right)\cdots ra_{r,0}^{\left(\lambda\right)}\left(r,x\right)\nonumber \\
 & =-r\sum_{i=0}^{N-r}\left(-1\right)^{i}\left(N\right)_{i}a_{r-1,0}^{\left(\lambda\right)}\left(N-i,x\right)+\left(-1\right)^{N-r+1}\left(N\right)_{N-r+1}\left(-1\right)^{r}r!\nonumber \\
 & =-r\sum_{i=0}^{N-r+1}\left(-1\right)^{i}\left(N\right)_{i}a_{r-1,0}^{\left(\lambda\right)}\left(N-i,x\right).\nonumber
\end{align}

Let $r=2$ in (\ref{eq:44}). Then, we have
\begin{align}
a_{2,0}^{\left(\lambda\right)}\left(N+1,x\right) & =-2\sum_{i=0}^{N-1}\left(-1\right)^{i}\left(N\right)_{i}a_{1,0}^{\left(\lambda\right)}\left(N-i,x\right)\label{eq:45}\\
 & =2\left(-1\right)^{N+1}\sum_{i=0}^{N+1}\left(N\right)_{i}\left(N-i-1\right)!\nonumber \\
 & =2\left(-1\right)^{N+1}N!\sum_{i=0}^{N-1}\frac{1}{N-i}\nonumber \\
 & =2\left(-1\right)^{N+1}N!\sum_{i=1}^{N}\frac{1}{i}\nonumber \\
 & =2\left(-1\right)^{N+1}N!H_{N,1}.\nonumber
\end{align}

For $r=3$ in (\ref{eq:44}), by (\ref{eq:45}), we get
\begin{align}
a_{3,0}^{\left(\lambda\right)}\left(N+1,x\right) & =-3\sum_{i=0}^{N-2}\left(-1\right)^{i}\left(N\right)_{i}a_{2,0}^{\left(\lambda\right)}\left(N-i,x\right)\label{eq:46}\\
 & =-3\sum_{i=0}^{N-2}\left(-1\right)^{i}\left(N\right)_{i}2\left(-1\right)^{N-i}\left(N-i-1\right)!H_{N-i-1,1}\nonumber \\
 & =3!\left(-1\right)^{N+1}N!\sum_{i=0}^{N-2}\frac{H_{N-i-1,1}}{N-i}\nonumber \\
 & =3!\left(-1\right)^{N+1}N!H_{N,2}.\nonumber
\end{align}

From (\ref{eq:46}), by $r=4$ in (\ref{eq:44}), we note that
\begin{align}
a_{4,0}^{\left(\lambda\right)}\left(N+1,x\right) & =-4\sum_{i=0}^{N-3}\left(-1\right)^{i}\left(N\right)_{i}a_{3,0}^{\left(\lambda\right)}\left(N-i,x\right)\label{eq:47}\\
 & =\left(-1\right)^{N+1}4!N!\sum_{i=0}^{N-3}\frac{H_{N-i-1,2}}{N-i}\nonumber \\
 & =\left(-1\right)^{N+1}4!N!H_{N,3}.\nonumber
\end{align}

Continuing this process, we get
\begin{equation}
a_{r,0}^{\left(\lambda\right)}\left(N+1,x\right)=\left(-1\right)^{N+1}N!r!H_{N,r-1},\quad\left(2\le r\le N\right).\label{eq:48}
\end{equation}

Let $2\le r\le N$, $1\le i\le r-1$. Then, by (\ref{eq:27}), we
get
\begin{align}
 & a_{i,r-i}^{\left(\lambda\right)}\left(N+1,x\right)\label{eq:49}\\
 & =\left(x+i-r+1\right)a_{i,r-i-1}^{\left(\lambda\right)}\left(N,x\right)-ia_{i-1,r-i}^{\left(\lambda\right)}\left(N,x\right)-Na_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)\nonumber \\
 & =\left(x+i-r+1\right)\left\{ a_{i,r-i-1}^{\left(\lambda\right)}\left(N,x\right)-Na_{i,r-i-1}^{\left(\lambda\right)}\left(N-1,x\right)\right\} \nonumber \\
 & \relphantom =-i\left\{ a_{i-1,r-i}^{\left(\lambda\right)}\left(N,x\right)-Na_{i-1,r-i}^{\left(\lambda\right)}\left(N-1,x\right)\right\} \nonumber \\
 & \relphantom =+\left(-1\right)^{2}N\left(N-1\right)a_{i,r-i}^{\left(\lambda\right)}\left(N-1,x\right)\nonumber \\
 & \vdots\nonumber \\
 & =\left(x+i-r+1\right)\sum_{s=0}^{N-r}\left(-1\right)^{s}\left(N\right)_{s}a_{i,r-i-1}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber\\
 &\relphantom{=}-i\sum_{s=0}^{N-r}\left(-1\right)^{s}\left(N\right)_{s}a_{i-1,r-i}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber \\
 & \relphantom =+\left(-1\right)^{N-r+1}\left(N\right)_{N-r+1}a_{i,r-i}^{\left(\lambda\right)}\left(r,x\right)\nonumber \\
 & =\left(x+i-r+1\right)\sum_{s=0}^{N-r}\left(-1\right)^{s}\left(N\right)_{s}a_{i,r-i-1}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber\\
 &\relphantom{=}-i\sum_{s=0}^{N-r}\left(-1\right)^{s}\left(N\right)_{s}a_{i-1,r-i}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber \\
 & \relphantom =+\left(-1\right)^{N-r+1}\left(N\right)_{N-r+1}\left(-1\right)^{i}i!S_{1,i-1}\left(r-i\right)\left(x\right)_{r-i}.\nonumber
\end{align}

Let $r=2$ in (\ref{eq:49}). Then $i=1$. From (\ref{eq:49}), we
note that
\begin{align}
 & a_{1,1}^{\left(\lambda\right)}\left(N+1,x\right)\label{eq:50}\\
 & =x\sum_{s=0}^{N-2}\left(-1\right)^{s}\left(N\right)_{s}a_{1,0}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber\\
 &\relphantom{=}-\sum_{s=0}^{N-2}\left(-1\right)^{s}\left(N\right)_{s}a_{0,1}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber \\
 & \relphantom =+\left(-1\right)^{N-1}N!\left(-1\right)S_{1,0}\left(1\right)x\nonumber \\
 & =x\sum_{s=0}^{N-2}\left(-1\right)^{s}\left(N\right)_{s}\left(-1\right)^{N-s}\left(N-s-1\right)!\nonumber\\
 &\relphantom{=}-\sum_{s=0}^{N-2}\left(-1\right)^{s}\left(N\right)_{s}\left(-1\right)^{N-s-1}\left(N-s-1\right)!x\nonumber \\
 & \relphantom =+2\left(-1\right)^{N}N!x\nonumber \\
 & =2x\left(-1\right)^{N}\sum_{s=0}^{N-2}\left(N\right)_{s}\left(N-s-1\right)!+2\left(-1\right)^{N}N!x\nonumber \\
 & =2x\left(-1\right)^{N}N!\sum_{s=0}^{N-2}\frac{1}{N-s}+2\left(-1\right)^{N}N!x\nonumber \\
 & =2\left(-1\right)^{N}N!x\left\{ \frac{1}{N}+\frac{1}{N-1}+\cdots+\frac{1}{2}+\frac{1}{1}\right\} \nonumber \\
 & =2\left(-1\right)^{N}N!xH_{N}.\nonumber
\end{align}

Let $r=3$. Then $1\le i\le2.$ From (\ref{eq:49}), we note that

\begin{align}
a_{i,3-i}^{\left(\lambda\right)}\left(N+1,x\right) & =\left(x+i-2\right)\sum_{s=0}^{N-3}\left(-1\right)^{s}\left(N\right)_{s}a_{i,2-i}^{\left(\lambda\right)}\left(N-s,x\right)\label{eq:51}\\
 & \relphantom{=}-i\sum_{s=0}^{N-3}\left(-1\right)^{s}\left(N\right)_{s}a_{i-1,3-i}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber \\
 & \relphantom{=}+\left(-1\right)^{N-2}\left(N\right)_{N-2}\left(-1\right)^{i}i!S_{1,i-1}\left(3-i\right)\left(x\right)_{3-i}.\nonumber
\end{align}

For $i=1$ in (\ref{eq:51}), we have
\begin{align}
 & a_{1,2}^{\left(\lambda\right)}\left(N+1,x\right)\label{eq:52}\\
 & =\left(x-1\right)\sum_{s=0}^{N-3}\left(-1\right)^{s}\left(N\right)_{s}a_{1,1}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber \\
 & \relphantom{=}-\sum_{s=0}^{N-3}\left(-1\right)^{s}\left(N\right)_{s}a_{0,2}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber \\
 & \relphantom{=}+\left(-1\right)^{N-2}\left(N\right)_{N-2}\left(-1\right)S_{1,0}\left(2\right)\left(x\right)_{2}\nonumber \\
 & =\left(x-1\right)\sum_{s=0}^{N-3}\left(-1\right)^{s}\left(N\right)_{s}2\left(-1\right)^{N-s-1}\left(N-s-1\right)!H_{N-s-1,1}x\nonumber \\
 & \relphantom{=}-\sum_{s=0}^{N-3}\left(-1\right)^{s}\left(N\right)_{s}\left(x\right)_{2}\left(-1\right)^{N-s-2}\left(N-s-1\right)!H_{N-s-1,1}\nonumber \\
 & \relphantom{=}+3\left(-1\right)^{N-1}\left(N\right)_{N-2}\left(x\right)_{2}\nonumber \\
 & =3\left(x\right)_{2}\left(-1\right)^{N-1}N!\sum_{s=0}^{N-3}\frac{H_{N-s-1,1}}{N-s}+3\left(x\right)_{2}\left(-1\right)^{N-1}N!\frac{1}{2}\nonumber \\
 & =3\left(x\right)_{2}\left(-1\right)^{N-1}N!\left\{ \frac{H_{N-1,1}}{N}+\frac{H_{N-2,1}}{N-1}+\cdots+\frac{H_{2,1}}{3}+\frac{H_{1,1}}{2}\right\} \nonumber \\
 & =3\left(x\right)_{2}\left(-1\right)^{N-1}N!H_{N,2}.\nonumber
\end{align}

Let $i=2$ in (\ref{eq:51}). From (\ref{eq:52}), we note that

\begin{align*}
a_{2,1}^{\left(\lambda\right)}\left(N+1,x\right) & =x\sum_{s=0}^{N-3}\left(-1\right)^{s}\left(N\right)_{s}a_{2,0}^{\left(\lambda\right)}\left(N-s,x\right)\\
&\relphantom{=}-2\sum_{s=0}^{N-3}\left(-1\right)^{s}\left(N\right)_{s}a_{1,1}^{\left(\lambda\right)}\left(N-s,x\right)\\
 & \relphantom =+\left(-1\right)^{N-2}\left(N\right)_{N-2}\left(-1\right)^{2}2!S_{1,1}\left(1\right)x\\
 & =x\sum_{s=0}^{N-3}\left(-1\right)^{s}\left(N\right)_{s}2\left(-1\right)^{N-s}\left(N-s-1\right)!H_{N-s-1,1}\\
 & \relphantom =-2\sum_{s=0}^{N-3}\left(-1\right)^{s}\left(N\right)_{s}2\left(-1\right)^{N-s-1}\left(N-s-1\right)!H_{N-s-1,1}x\\
 & \relphantom =+\left(-1\right)^{N-2}\left(N\right)_{N-2}\left(-1\right)^{2}3!x\\
 & =6x\left(-1\right)^{N}N!\sum_{s=0}^{N-3}\frac{H_{N-s-1,1}}{N-s}+6x\left(-1\right)^{N}\left(N\right)_{N-2}\\
 & =6x\left(-1\right)^{N}N!\left\{ \frac{H_{N-1,1}}{N}+\cdots+\frac{H_{2,1}}{3}+\frac{H_{1,1}}{2}\right\} \\
 & =6x\left(-1\right)^{N}N!H_{N,2}.
\end{align*}

Therefore, we obtain the following theorem.
\begin{thm}
\label{thm:2} Let $1\le r\le N+1$, $0\le i\le r$. Then we have
\[
a_{i,r-i}^{\left(\lambda\right)}\left(N+1,x\right)=\left(-1\right)^{N+1+i-r}i!S_{1,i-1}\left(r-i\right)N!H_{N,r-1}\left(x\right)_{r-i}.
\]
\end{thm}
\begin{proof}
We showed that it is true for $r=1$ and $r=N+1$. Assume that $2\le r\le N$.
If $i=0$ or $i=r$, then it is also true. So we prove the assertion
by inductino on $r$ when $2\le r\le N$, $1\le i\le r-1$. For $r=2$,
$i=1$, we showed that
\[
a_{1,1}^{\left(\lambda\right)}\left(N+1,x\right)=2\left(-1\right)^{N}N!H_{N,1}x.
\]

Assume now that it is true for $r-1\left(3\le r\le N\right)$.

From \ref{eq:49}, we note that
\begin{align}
&\relphantom{=}a_{i,r-i}^{\left(\lambda\right)}\left(N+1,x\right)\label{eq:53} \\
& =\left(x+i-r+1\right)\sum_{s=0}^{N-r}\left(-1\right)^{s}\left(N\right)_{s}a_{i,r-i-1}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber\\
  & \relphantom =-i\sum_{s=0}^{N-r}\left(-1\right)^{s}\left(N\right)_{s}a_{i-1,r-i}^{\left(\lambda\right)}\left(N-s,x\right)\nonumber\\
 & \relphantom =+\left(-1\right)^{N-r+1}\left(N\right)_{N-r+1}\left(-1\right)^{i}i!S_{1,i-1}\left(r-i\right)\left(x\right)_{r-i}\nonumber \\
 & =\left(x+i-r+1\right)\sum_{s=0}^{N-r}\left(-1\right)^{s}\left(N\right)_{s}\left(-1\right)^{N-s-1+i-r}i!\nonumber\\
 &\relphantom {=}\times S_{1,i-1}\left(r-1-i\right)\left(N-s-1\right)!H_{N-s-1,r-2}\left(x\right)_{r-1-i}\nonumber \\
 & \relphantom {=}-i\sum_{s=0}^{N-r}\left(-1\right)^{s}\left(N\right)_{s}\left(-1\right)^{N-s+i-r}\left(i-1\right)!S_{1,i-2}\left(r-i\right)\nonumber\\
 &\relphantom{=}\times\left(N-s-1\right)!H_{N-s-1,r-2}\left(x\right)_{r-i}\nonumber\\
 & \relphantom {=}+\left(-1\right)^{N-r+1}\left(N\right)_{N-r+1}\left(-1\right)^{i}i!S_{1,i-1}\left(r-i\right)\left(x\right)_{r-i}\nonumber\\
 & =\left(-1\right)^{N+1+i-r}i!N!\left(x\right)_{r-i}\nonumber\\
 &\relphantom{=}\times\left\{ \left(S_{1,i-1}\left(r-1-i\right)+S_{1,i-2}\left(r-i\right)\right)\sum_{s=0}^{N-r}\frac{H_{N-s-1,r-2}}{N-s}+\frac{S_{1,i-1}\left(r-i\right)}{\left(r-1\right)!}\right\} .\nonumber
\end{align}
 By Lemma \ref{lem:1} and (\ref{eq:53}), we get
\begin{align}
a_{i,r-i}^{\left(\lambda\right)}\left(N+1,x\right) & =\left(-1\right)^{N+1+i-r}i!S_{1,i-1}\left(r-i\right)N!\left(x\right)_{r-i}\label{eq:54}\\
&\relphantom{=}\times\left\{ \frac{H_{N-1,r-2}}{N}+\cdots+\frac{H_{r-1,r-2}}{r}+\frac{H_{r-2,r-2}}{r-1}\right\} \nonumber\\
 & =\left(-1\right)^{N+1+i-r}i!S_{1,i-1}\left(r-i\right)N!H_{N,r-1}\left(x\right)_{r-i}.\qedhere\nonumber
\end{align}

\end{proof}
Therefore, we obtain the following theorem.
\begin{thm}
\label{thm:3} The linear differential equations
\begin{align*}
&\relphantom{=}F^{\left(N\right)}\\
&=\lambda^{N}\left(1+\lambda t\right)^{-N}\\
&\relphantom{=}\times\left(\sum_{r=1}^{N}\sum_{i=0}^{r}a_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-i}\left(\lambda+\log\left(1+\lambda t\right)\right)^{-\left(r-i\right)}\right)F
\end{align*}
has a solution
\[
F=F\left(t;x,\lambda\right)=2\lambda\left(2\lambda+\log\left(1+\lambda t\right)\right)^{-1}\left(1+\lambda^{-1}\log\left(1+\lambda t\right)\right)^{x},
\]
where, for $1\le r\le N$, $0\le i\le r$,
\[
a_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)=\left(-1\right)^{N+i-r}i!S_{1,i-1}\left(r-i\right)\left(N-1\right)!H_{N-1,r-1}\left(x\right)_{r-i}.
\]

\end{thm}
Recall that the $\lambda$-Changhee polynomials, $\Ch_{n,\lambda}\left(x\right)$,
$\left(n\ge0\right)$, are given by the generating function
\begin{align}
F & =F\left(t;x,\lambda\right)\label{eq:55}\\
 & =\frac{2\lambda}{2\lambda+\log\left(1+\lambda t\right)}\left(1+\frac{1}{\lambda}\log\left(1+\lambda t\right)\right)^{x}\nonumber \\
 & =\sum_{n=0}^{\infty}\Ch_{n,\lambda}\left(x\right)\frac{t^{n}}{n!}.\nonumber
\end{align}

Thus, by (\ref{eq:55}), we get
\begin{align}
F^{\left(N\right)} & =\left(\frac{d}{dt}\right)^{N}F\left(t;x,\lambda\right)\label{eq:56}\\
 & =\sum_{k=0}^{\infty}\Ch_{k+N,\lambda}\left(x\right)\frac{t^{k}}{k!}.\nonumber
\end{align}

On the other hand, by Theorem \ref{thm:3}, we get
\begin{align}
 & F^{\left(N\right)}\label{eq:57}\\
 & =\lambda^{N}\sum_{l=0}^{\infty}\left(-1\right)^{l}\binom{N+l-1}{l}\lambda^{l}t^{l}\sum_{r=1}^{N}\sum_{i=0}^{r}a_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)\sum_{m=0}^{\infty}\left(-1\right)^{m}\binom{i+m-1}{m}\nonumber \\
 & =\left(2\lambda\right)^{-i-m}\left(\log\left(1+\lambda t\right)\right)^{m}\sum_{n=0}^{\infty}\left(-1\right)^{n}\binom{r+n-i-1}{n}\lambda^{-\left(r-i\right)-n}\left(\log\left(1+\lambda t\right)\right)^{n}\nonumber \\
 & \relphantom =\times\sum_{s=0}^{\infty}\Ch_{s,\lambda}\left(x\right)\frac{t^{s}}{s!}\nonumber \\
 & =\lambda^{N}\sum_{l=0}^{\infty}\left(-1\right)^{l}\left(N+l-1\right)_{l}\lambda^{l}\frac{t^{l}}{l!}\nonumber\\
 &\relphantom{=}\times\sum_{r=1}^{N}\sum_{i=0}^{r}a_{i,N-i}^{\left(\lambda\right)}\left(N,x\right)\sum_{m=0}^{\infty}\left(-1\right)^{m}\left(i+m-1\right)_{m}\left(2\lambda\right)^{-i-m}\nonumber \\
  &\relphantom{=}\times\sum_{e=0}^{\infty}S_{1}\left(e+m,m\right)\lambda^{e+m}\frac{t^{e+m}}{\left(e+m\right)!}\sum_{n=0}^{\infty}\left(-1\right)^{n}\left(r+n-i-1\right)_{n}\nonumber\\
 &\relphantom{=}\times\lambda^{i-r-n}\sum_{f=0}^{\infty}S_{1}\left(f+n,n\right)\nonumber \\
 & \relphantom =\times\lambda^{f+n}\frac{t^{f+n}}{\left(f+n\right)!}\sum_{s=0}^{\infty}\Ch_{s,\lambda}\left(x\right)\frac{t^{s}}{s!}\nonumber \\
 & =\lambda^{N}\sum_{r=1}^{N}\sum_{i=0}^{r}a_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)\sum_{m=0}^{\infty}\left(-1\right)^{m}\left(i+m-1\right)_{m}\left(2\lambda\right)^{-i-m}\lambda^{m}\frac{t^{m}}{m!}\nonumber \\
 & \relphantom =\times\sum_{n=0}^{\infty}\left(-1\right)^{n}\left(r+n-i-1\right)_{n}\lambda^{i-r-n}\lambda^{n}\frac{t^{n}}{n!}\sum_{l=0}^{\infty}\left(-1\right)^{l}\left(N+l-1\right)_{l}\lambda^{l}\frac{t^{l}}{l!}\nonumber \\
 & \relphantom =\times\sum_{e=0}^{\infty}S_{1}\left(e+m,m\right)\lambda^{e}\frac{e!m!}{\left(e+m\right)!}\frac{t^{e}}{e!}\sum_{f=0}^{\infty}S_{1}\left(f+n,n\right)\lambda^{f}\frac{f!n!}{\left(f+n\right)!}\frac{t^{f}}{f!}\nonumber \\
 & \relphantom =\times\sum_{s=0}^{\infty}\Ch_{s,\lambda}\left(x\right)\frac{t^{s}}{s!}\nonumber\\
 & =\lambda^{N}\sum_{r=1}^{N}\sum_{i=0}^{r}a_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)\sum_{m=0}^{\infty}\left(-1\right)^{m}\nonumber\\
 &\relphantom{=}\times\left(i+m-1\right)_{m}\left(2\lambda\right)^{-i-m}\lambda^{m}\frac{t^{m}}{m!}\sum_{n=0}^{\infty}\left(-1\right)^{n}\left(r+n-i-1\right)_{n}\nonumber \\
 & \relphantom =\times\lambda^{i-r-n}\lambda^{n}\frac{t^{n}}{n!}\sum_{a=0}^{\infty}\left(\sum_{l+e+f+s=a}\left(-1\right)^{l}\lambda^{a-s}\frac{\binom{a}{l,e,f,s}}{\binom{e+m}{m}\binom{f+n}{n}}\left(N+l-1\right)_{l}\right.\nonumber\\
 &\relphantom{=}\left.S_{1}\left(e+m,m\right)S_{1}\left(f+n,n\right)\Ch_{s,\lambda}\left(x\right)\right)\frac{t^{a}}{a!}\nonumber \\
 & =\lambda^{N}\sum_{r=1}^{N}\sum_{i=0}^{r}a_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)\lambda^{-r}2^{-i}\sum_{k=0}^{\infty}\left(\sum_{m+n+a=k}\binom{k}{m,n,a}\left(-\frac{1}{2}\right)^{m}\left(-1\right)^{n}\left(i+m-1\right)_{m}\right.\nonumber \\
 & \relphantom =\times\left(r+n-i-1\right)_{n}\sum_{l+e+f+s=a}\left(-1\right)^{l}\lambda^{a-s}\frac{\binom{a}{l,e,f,s,}}{\binom{e+m}{m}\binom{f+n}{n}}\nonumber\\
 &\relphantom{=}\times \left(N+l-1\right)_{l}S_{1}\left(e+m,m\right)\left. S_{1}\left(f+n,n\right)\Ch_{s,\lambda}\left(x\right)\right)\frac{t^{k}}{k!}\nonumber \\
 & =\sum_{k=0}^{\infty}\left\{ \lambda^{N}\sum_{r=1}^{N}\sum_{i=0}^{r}a_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)\lambda^{-r}2^{-i}\right.\nonumber \\
 & \relphantom =\times\sum_{m+n+a=k}\binom{k}{m,n,a}\left(-\frac{1}{2}\right)^{m}\left(-1\right)^{n}\left(i+m-1\right)_{m}\left(r+n-i-1\right)_{n}\nonumber \\
 & \relphantom =\times\sum_{l+e+f+s=a}\left(-1\right)^{l}\lambda^{a-s}\frac{\binom{a}{l,e,f,s}}{\binom{e+m}{m}\binom{f+n}{n}}\left(N+l-1\right)_{l}\nonumber \\
 & \relphantom =\left.\times S_{1}\left(e+m,m\right)S_{1}\left(f+n,n\right)\Ch_{s,\lambda}\left(x\right)\right\} \frac{t^{k}}{k!}.\nonumber
\end{align}

Therefore, by comparing the coefficients on both sides of (\ref{eq:56})
and (\ref{eq:57}), we obtain the following theorem.
\begin{thm}
\label{thm:4} For $N\in\mathbb{N},$ and $k\in\mathbb{N}\cup\left\{ 0\right\} $,
we have
\begin{align*}
 & \Ch_{k+N,\lambda}\left(x\right)\\
 & =\lambda^{N}\sum_{r=1}^{N}\sum_{i=0}^{r}a_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)\lambda^{-r}2^{-i}\sum_{m+n+a=k}\binom{k}{m,n,a}\\
 & \relphantom =\times\left(-\frac{1}{2}\right)^{m}\left(-1\right)^{n}\left(i+m-1\right)_{m}\left(r+n-i-1\right)_{n}\\
 & \relphantom =\times\sum_{l+e+f+s=a}\left(-1\right)^{l}\lambda^{a-s}\frac{\binom{a}{l,e,f,s}}{\binom{e+m}{m}\binom{f+n}{n}}\\
 & \relphantom =\times\left(N+l-1\right)_{l}S_{1}\left(e+m,m\right)S_{1}\left(f+n,n\right)\Ch_{s,\lambda}\left(x\right),
\end{align*}
where, for $1\le r\le N$, $0\le i\le r$,
\[
a_{i,r-i}^{\left(\lambda\right)}\left(N,x\right)=\left(-1\right)^{N+i-r}i!S_{1,i-1}\left(r-i\right)\left(N-1\right)!H_{N-1,r-1}\left(x\right)_{r-i}.
\]

\end{thm}
\bibliographystyle{amsplain}
\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}

\end{document}